Question 2 (10 points) Z is a standard normal random variable. The P(z 2 2.11) equals 0.4821 0.9821 0.5 0.0174
Balthus Corp. reports the following components of stockholders equity on December 31, 2016: Problem 13-2B Cash dividends, treasury stock, and statement of retained earnings C3 P2 P3 Common stock-$1 par value. 320,000 shares authorized, 200,000 shares issued and outstanding Paid-in capital in excess of par value, common stock Retained earnings . . 200,000 1,400,000 2,160.000 It completed the following transactions related to stockholders equity in year 2017: Jan. 10 Mar. 2 Purchased 40,000 shares of its own stock at $12 cash per share. Directors declared a $1.50 per share cash dividend payable on March 31 to the March 15 stock- holders of record. Mar. 31 Paid the dividend declared on March 2 Nov. 11 Sold 24,000 of its treasury shares at $13 cash per share. Nov. 25 Sold 16,000 of its treasury shares at $9.50 cash per share. Dec. 1 Directors declared a $2.50 per share cash dividend payable on January 2 to the December 10 stockholders of record. Closed the $1,072,000 credit balance (from net income) in the Income Summary account to Retained Earnings. Dec. 31 Required 1. Prepare journal entries to record each of these transactions for 2017 2. Prepare a statement of retained earnings for the year ended December 31.2017. 3. Prepare the stockholders equity section of the companys balance sheet as of December 31, 2017. Check (2) Retained earnings Dec 31, 2017, $2,476,000
Answer #1

Solution : ( 2 )

P(z > 2.11) 1-P(z < 2.11) = 1-0.98257= 0.01743 0.0174

Hence, P (z > 2.11) = 0.0174